Find the equation of a line passing through the point (- 2, - 4) and parallel to the line 2x + 3y = 24.
Answer by mananth(16946)

Find the equation of a line passing through the point (- 2, - 4) and parallel to the line 2x + 3y = 24.
....
(- 2, - 4) and parallel to the line
2x + 3y = 24
3y=-2x+24
/3
y=(-2x/3)+8
...
slope m = (-2/3)
...
Parallel line will have same slope = (-2/3)
find b
y= mx+b
(-2,-4) m=(-2/3)
plug these values
-4=(-2/3)*-2 +b
-4=4/3+b
b=-4-4/3
b=(-16/3)
...
equation is
y=(-2/3)x-16/3
...
m.ananth@hotmail.ca